# Neural Collapse-Inspired Multi-Label Federated Learning under Label-Distribution Skew

### v2

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy. However, the performance of deep learning often deteriorates in FL due to decentralized and heterogeneous data. This challenge is further amplified in multi-label scenarios, where data exhibit complex characteristics such as label co-occurrence, inter-label dependency, and discrepancies between local and global label relationships. While most existing FL research primarily focuses on single-label classification, many real-world applications, particularly in domains such as medical imaging, often involve multi-label settings. In this paper, we address this important yet underexplored scenario in FL, where clients hold multi-label data with skewed label distributions. Neural Collapse (NC) describes a geometric structure in the latent feature space where features of each class collapse to their class mean with vanishing intra-class variance, and the class means form a maximally separated configuration. Motivated by this theory, we propose a method to align feature distributions across clients and to learn high-quality, well-clustered representations. To make the NC-structure applicable to multi-label settings, where image-level features may contain multiple semantic concepts, we introduce a feature disentanglement module that extracts semantically specific features. The clustering of these disentangled class-wise features is guided by a predefined shared NC structure, which mitigates potential conflicts between client models due to diverse local data distributions. In addition, we design regularisation losses to encourage compact clustering in the latent feature space. Experiments conducted on four benchmark datasets across eight diverse settings demonstrate that our approach outperforms existing methods, validating its effectiveness in this challenging FL scenario.
Introduction
Figure 1: Within an FL system, clients may present distinct data distributions and label relationships. For example, Client 1 primarily serves patients with sun exposure-related skin lesions, whereas Client 2 mainly handles cases involving melanoma and vascular lesions. The proposed method aims to address multi-label, label-skewed FL scenarios by disentangling class-specific features and promoting structured feature clustering in the latent feature space.
@@ Introduction
Federated Learning (FL) has emerged as a promising framework for analysing sensitive data, particularly in domains like healthcare imaging, where privacy and legal constraints prevent direct data sharing. However, most existing FL approaches are limited to basic multi-class classification tasks, failing to account for more realistic scenarios where multiple labels frequently co-occur within a single sample. For instance, many diseases are interrelated, and patients often present with multiple conditions or symptoms simultaneously, rather than a single isolated issue. This multi-label nature of real-world data is often overlooked. Another core challenge in FL is data heterogeneity across clients, as each client (e.g., hospital) may possess a uniquely skewed data distribution. This variation arises from differences in geography, population demographics, medical facilities, and clinical expertise. As a result, clients may exhibit label imbalance, including majority, minority, or even missing (vacant) classes. Under multi-label scenarios, such heterogeneity becomes even more pronounced, making collaboration across clients significantly more difficult. Figure 1 illustrates this scenario, demonstrating that hospitals may hold datasets containing various sets of dermatological conditions.
@@
In this paper, we explore the problem of multi-label FL under label-skewed conditions, where each client’s local data is both imbalanced and divergent from the global distribution. Without directly sharing local data, the goal is to collaboratively train a global model that performs well across all target classes, despite the presence of class imbalance and missing labels on certain clients. To tackle the unique challenges posed by multi-label, label-skewed FL, we propose a novel method called Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML). Our approach introduces a representation learning framework for multi-label FL, inspired by Neural Collapse (NC) theory (Papyan et al.,, 2020), to enhance feature alignment across clients with heterogeneous label distributions. Specifically, we first encode each input image using a standard feature extractor, producing an image-level representation potentially containing multiple semantic concepts. We then leverage attention mechanisms to extract class-specific representations from this image-level representation, generating disentangled, class-wise embeddings. This effectively decomposes the complex multi-label problem into multiple single-label problems. Next, motivated by the NC phenomenon, we incorporate a predefined Equiangular Tight Frame (ETF) structure to promote consistent and discriminative feature alignment across clients, even under non-IID data distributions. The ETF structure facilitates class-independent clustering, encouraging local models to align with the global class geometry rather than overfitting to their local data distributions. Additionally, we introduce two complementary regularisation terms to enhance feature clustering in the latent feature space. A rejection loss filters out noisy signals from negative features, and a contrastive loss encourages compact intra-class clustering and clear inter-class separation. The key contributions of this paper are as follows:
@@
  • We explore the important yet underexplored problem of multi-label FL under label-skewed conditions. To this end, we propose FedNCAlign-ML, a novel method inspired by Neural Collapse theory.
  • FedNCAlign-ML effectively adapts established representation learning techniques from the multi-class to the multi-label setting, and tackles the challenges introduced by heterogeneous local data distributions, label co-occurrence, and inter-label dependency.
  • To handle the presence of multiple semantic concepts within multi-label samples, we introduce an attention-based feature disentanglement block that extracts class-specific representations from image-level features. Additionally, we incorporate two regularisation terms, one to suppress noisy negative features and another to promote compact class-wise clustering.
  • We evaluate the proposed FedNCAlign-ML method on four benchmark datasets under eight diverse settings, comparing it to state-of-the-art FL approaches. The results show that FedNCAlign-ML consistently outperforms existing methods, achieving improvements of up to 2.87% in class-wise AUC and 7.29% in class-wise F1 score.
Related Work
Heterogeneous Federated Learning
@@
The foundational FL algorithm, FedAvg (McMahan et al.,, 2017), performs iterative weighted averaging of locally trained model parameters. This process involves local training on client devices, aggregation at a central server, and broadcasting the updated global model back to clients. While FedAvg performs well under independently and identically distributed (IID) data, real-world scenarios often involve non-IID data distributions. These distributions are commonly observed as quantity skew, label distribution skew, and feature distribution skew, all of which present significant challenges for FL. In the context of multi-class classification, numerous methods have been proposed to address such data heterogeneity within FL settings. For instance, FedProx (Li et al.,, 2020) adds an L2 regularisation term to constrain local updates and encourage proximity to the global model. FedCurv (Shoham et al.,, 2019) uses Elastic Weight Consolidation (Kirkpatrick et al.,, 2017) to preserve important local parameters to reduce divergence across clients. SCAFFOLD (Karimireddy et al.,, 2020) introduces control variates to correct client drift. SphereFed (Dong et al.,, 2022) projects features onto a hypersphere using a fixed classifier to mitigate latent feature distribution discrepancies. FedVLS (Guo et al.,, 2025) applies global model distillation to alleviate forgetting of missing classes and regularisation to amplify minority class losses, thereby improving performance on underrepresented classes. Although these approaches have demonstrated effectiveness in the multi-class setting, their performance tends to deteriorate when directly applied to the more complex multi-label scenario, where each sample may be associated with multiple co-occurring labels. In this work, we focus on addressing the challenges of multi-label FL under quantity-skewed and label-skewed data distribution.
Multi-Label Federated Learning
@@
Although multi-label classification is a fundamental task, it remains less explored than multi-class classification in image processing. In the context of FL, the multi-label setting introduces additional and distinct challenges compared to the multi-class scenario, including label co-occurrence, inter-label dependency, and divergence between local and global label relationships. FedLGT (Liu et al.,, 2024) addresses these challenges by focusing on modelling label correlations across clients. Building upon C-Tran (Lanchantin et al.,, 2021), it leverages frozen CLIP (Radford et al.,, 2021) text embeddings to maintain consistent label relationships and reduce divergence in label dependencies among clients. In contrast, our approach focuses on understanding the structure of the latent feature space and explicitly controlling and optimising it to enhance robustness under challenging label-skewed FL conditions.
@@
Figure 2: Framework for the proposed FedNCAlign‑ML method for addressing multi‑label, label‑skewed FL. First, an attention‑based Label‑Aware Disentanglement Module (LADM) extracts label‑specific features from image‑level representations. Then, a predefined ETF matrix serves as both the shared classifier and the source of class-wise query embeddings, enabling consistent local training across clients. Additionally, two regularisation terms are introduced to suppress noisy negative features and encourage compact intra-class clustering in the latent feature space.
Neural Collapse-Inspired Methods
@@
Recent research has shown that during the terminal phase of training (TPT), a phenomenon known as Neural Collapse (NC) emerges (Papyan et al.,, 2020). TPT refers to the stage when overparameterized deep networks are trained on balanced multi-class classification datasets, and the model’s performance plateaus while the training loss remains non-zero and continues to decrease. NC describes a distinctive geometric structure in the latent feature space: final-layer features collapse to their corresponding within-class means, and these class-wise prototypes align to form the vertices of a simplex Equiangular Tight Frame (ETF). As this simple yet elegant mathematical structure provides an optimal linear separable state for classification, NC-inspired classifiers have been explored in various domains, including federated learning (Li et al., 2023b, ), incremental learning (Yang et al., 2023b, ; Wei et al.,, 2025), few-shot learning (Yang et al., 2023b, ), and long-tail recognition (Gao et al.,, 2024). Despite growing interest, most existing work has focused on multi-class classification, while the equally important challenge of multi-label classification remains underexplored. To address this gap, Li et al., 2023a investigates the presence of NC in multi-label settings. They find that network training on a balanced multi-label dataset also exhibits NC but in a generalised form. In particular, features corresponding to single-label instances still conform to the Simplex ETF structure, while features associated with multi-label instances tend to approximate scaled averages of the single-label prototypes. MLC-NC (Tao et al.,, 2025) uses NC to process the long-tailed multi-label classification task in a centralised setting. However, it does not consider the missing class condition, which frequently arises in label-skewed FL scenarios across individual clients. To the best of our knowledge, our paper is the first to explore multi-label FL from an NC perspective. It opens the possibility to solve the heterogeneous multi-label FL challenge by fundamentally avoiding interference due to various local data distributions, label co-occurrence, and inter-label dependency.
Preliminaries
Problem Formulation
@@
We consider a multi-label, label-skewed FL setting with $K$ clients collaboratively training a shared global model. Each client $k$ holds a private dataset $\mathcal{D}_{k}=\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{N_{k}}$ of size $N_{k}$, where the input image $x_{i}^{k}\in\mathbb{R}^{H\times W\times 3}$ and the label $y_{i}^{k}\in\{0,1\}^{C}$ is a multi-hot vector over the $C$ classes. Let $\mathcal{C}=\{1,\dots,C\}$ denote the global class index set. Due to label skew, client $k$ may only observe a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients. The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing raw data among clients.
Neural Collapse (NC)
@@
In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of Neural Collapse.
@@
  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses: features of the same class concentrate at their class mean.
@@
  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
@@
  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, sharing the same simplex ETF geometry.
Proposed Method
The NC phenomenon suggests a promising structure for optimal feature clustering in the latent feature space under well-trained conditions. We leverage this geometry to shape the latent feature space in the challenging multi-label FL setting, where client data often deviates from the global distribution. An overview of our framework is shown in Figure 2.
Label-Aware Disentanglement Module (LADM)
Multi-label images may contain multiple object categories, making a single pooled representation suboptimal because it compresses multiple semantics into one vector. Inspired by recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label.
[…]

### v3

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy, yet it remains challenging as data distributions can be highly heterogeneous. These challenges are further amplified in multi-label scenarios, where data exhibit characteristics such as label co-occurrence, inter-label dependency, and discrepancies between local and global label relationships. While most existing FL studies focus on single-label classification, real-world applications, such as in medical imaging, involve multi-label data with highly skewed label distributions across clients. To address this important yet underexplored problem, we propose FedNCA-ML, a novel FL framework that aligns feature distributions across clients and learns discriminative, well-clustered representations inspired by Neural Collapse (NC) theory. NC describes an ideal latent-space geometry where each class’s features collapse to their mean, forming a maximally separated simplex. To extend this theory to multi-label settings, we introduce a feature disentanglement module that extracts class-specific representations. The clustering of these disentangled features is guided by a shared NC-inspired structure, mitigating conflicts among client models caused by heterogeneous local data. Furthermore, we design regularisation losses to encourage compact and consistent feature clustering in the latent space. Experiments on four benchmark datasets under eight FL settings demonstrate the effectiveness of the proposed method, achieving improvements of up to 3.92% in class-wise AUC and 4.93% in class-wise F1 score.
1 Introduction
Figure 1: (a): Local client data are heterogeneous, exhibiting distinct class imbalances. (b): Multi-label data further exacerbate this heterogeneity through diverse label correlations. (c): FedNCA-ML addresses multi-label label-skewed FL challenges by disentangling class-specific features and promoting Neural Collapse–inspired structured clustering in the latent space.
@@ 1 Introduction
Federated Learning (FL) has emerged as a promising framework for analysing sensitive data, particularly in domains such as medical imaging, where privacy regulations and legal constraints prohibit direct data sharing. However, most existing FL approaches focus on basic multi-class classification tasks, overlooking more realistic scenarios in which multiple labels often co-occur within a single sample. For instance, many diseases are interrelated, and patients frequently present with multiple concurrent conditions. A fundamental challenge in FL is data heterogeneity across clients. Each client (e.g., hospital) may have a uniquely skewed data distribution due to variations in geography, population demographics, medical facilities, and clinical expertise. Consequently, local datasets often exhibit label imbalance, including majority, minority, or even missing classes. In multi-label scenarios, such heterogeneity becomes even more pronounced. Within each client, certain labels may frequently co-occur, forming client-specific label correlations that differ across clients. These inconsistent label relationships further complicate collaborative model training, posing significant challenges to achieving global model convergence and generalization. Figure 1 illustrates this scenario, showing that different clients hold datasets with diverse data distributions and label dependencies.
@@
In this paper, we explore the problem of multi-label FL under label-skewed conditions, where each client’s local data is both imbalanced and diverges from the global distribution. Without directly sharing local data, the goal is to collaboratively train a global model that performs well across all target classes, despite class imbalance, missing labels on certain clients, and diverse label correlations across clients. To tackle these challenges, we propose a novel method called Federated Neural Collapse Alignment for Multi-Label Learning (FedNCA-ML). FedNCA-ML introduces a unified representation learning framework for multi-label FL, inspired by Neural Collapse (NC) theory (papyan2020prevalence), to achieve consistent and discriminative feature alignment across heterogeneous clients. Specifically, each input image is first encoded by a standard feature extractor to obtain an image-level representation containing multiple semantic concepts. An attention-based disentanglement module then derives class-specific representations, effectively decomposing the complex multi-label problem into multiple single-label subproblems. Guided by NC theory, we further integrate a predefined Equiangular Tight Frame (ETF) structure to enforce consistent and discriminative feature alignment across clients. The ETF structure facilitates class-independent clustering, encouraging local models to align with the global class geometry rather than overfitting to their local data distributions. In addition, two complementary regularisation terms enhance feature compactness and robustness. A rejection loss filters out noisy negative features, while a contrastive loss encourages compact intra-class clustering and clear inter-class separation. The key contributions of this paper are as follows:
@@
  • We explore the important yet underexplored problem of multi-label FL under label-skewed conditions, and propose FedNCA-ML, a novel NC-inspired framework for representation alignment in heterogeneous multi-label settings.
  • FedNCA-ML proposes an attention-based disentanglement module to extract class-specific representations, extending NC principles to multi-label learning and mitigating data heterogeneity and label dependency across clients.
  • Two regularization terms are designed to suppress noisy negative features and encourage compact class-wise compactness, improving global model generalization under label-skewed FL.
2 Related Work
2.1 Heterogeneous Federated Learning
@@
The classical FL algorithm FedAvg (mcmahan2017communication), which iteratively performs weighted averaging of locally trained model parameters, remains the foundation of most FL methods. Its process involves local training on client devices, aggregation at a central server, and broadcasting of the updated global model back to all clients. While FedAvg performs well under independently and identically distributed (IID) data, real-world applications often involve non-IID data distributions, which significantly hinder FL performance. Such heterogeneity typically arises from quantity skew, label distribution skew, and feature distribution skew. In the context of multi-class classification, many approaches have been proposed to alleviate these challenges in FL (li2020federated; shoham2019overcoming; karimireddy2020scaffold; dong2022spherefed; guo2025exploring; saha2025f). A detailed discussion of these works is provided in the Appendix A. However, their effectiveness degrades when directly applied to multi-label scenarios, where each sample may contain multiple co-occurring labels. In this work, we focus on addressing multi-label FL under quantity-skewed and label-skewed data distributions.
2.2 Multi-Label Federated Learning
@@
In FL, the multi-label setting poses unique challenges, including label co-occurrence, inter-label dependency, and divergence between local and global label relationships. FedMLP (sun2024fedmlp) addresses these issues by allowing each client to exchange both its local model and class-wise positive and negative prototypes with the server to perform pseudo-labeling. FedLGT (liu2024language) focuses on modelling label correlations across clients. Building on C-Tran (lanchantin2021general), it leverages frozen CLIP (radford2021learning) text embeddings to maintain consistent label relationships and reduce divergence in label dependencies among clients. In contrast, our approach focuses on structuring and optimizing the latent feature space, explicitly enhancing its organization to improve model robustness under label-skewed multi-label FL conditions.
@@
Figure 2: Overview of the proposed FedNCA-ML framework for multi-label label-skewed FL. Subfigure (a) shows the overall architecture, while Subfigures (b)–(d) illustrate the Label-Aware Disentanglement Module (LADM) and the regularization losses. The attention-based LADM extracts label-specific features from image-level features. A predefined ETF matrix acts as both the shared classifier and the source of class-wise query embeddings, ensuring consistent local training across clients. Two regularisation terms are further incorporated to suppress noisy negative features and promote compact intra-class clustering in the latent feature space.
2.3 Neural Collapse-Inspired Methods
@@
Recent studies have shown that during the terminal phase of training (TPT), a phenomenon known as Neural Collapse (NC) emerges (papyan2020prevalence). TPT refers to the stage in which overparameterized deep networks, trained on balanced multi-class classification datasets, reach performance saturation while the training loss continues to decrease. NC describes a distinctive geometric structure in the latent feature space: final-layer features collapse to their corresponding within-class means, and these class-wise prototypes align to form the vertices of a simplex Equiangular Tight Frame (ETF). This simple yet elegant geometric structure represents an optimally separable state for classification, motivating the development of NC-inspired classifiers across various tasks (li2023no; yang2023neural; wei2025compress; yang2023neural; gao2024distribution). Despite growing interest, most existing works focus on multi-class classification, leaving the equally important problem of multi-label classification underexplored. To bridge this gap, li2023neural investigated NC in multi-label settings and observed that models trained on balanced multi-label datasets exhibit a generalized form of NC. In particular, features corresponding to single-label instances still conform to the Simplex ETF structure, whereas features of multi-label samples approximate scaled averages of single-label prototypes. MLC-NC (tao2025mlc) further leverages NC to address long-tailed multi-label classification in a centralised setting. To the best of our knowledge, our paper is the first to explore multi-label FL from an NC perspective, opening new opportunities to address heterogeneous multi-label FL by mitigating interference caused by diverse local data distributions, label co-occurrence, and inter-label dependencies.
3 Preliminaries
3.1 Problem Formulation
@@
We consider a multi-label, label-skewed FL setting with $K$ clients collaboratively training a shared global model. Each client $k$ holds a private dataset $\mathcal{D}_{k}=\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{N_{k}}$ of size $N_{k}$, where the input image $x_{i}^{k}\in\mathbb{R}^{H\times W\times 3}$ and the label $y_{i}^{k}\in\{0,1\}^{C}$ is a multi-hot vector over the $C$ classes. Let $\mathcal{C}=\{1,\dots,C\}$ denote the global class index set. Due to label skew, client $k$ only observes a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients. The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing any local client data.
3.2 Neural Collapse (NC)
@@
In this section, we first introduce the structure of the simplex ETF, followed by a discussion of the key properties of NC.
@@
  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses, causing features of the same class to concentrate around their class mean.
@@
  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange themselves as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
@@
  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, exhibiting the same simplex ETF geometry.
4 Proposed Method
NC theory suggests a promising structure for optimal feature clustering in the latent space. Building on this insight, we propose FedNCA-ML, a framework that leverages NC geometry to organize and regularize feature representations in the multi-label FL setting. An overview of the proposed framework is presented in Figure 2. Our method employs an attention-based disentanglement module (Fig. 2b) to extract class-specific features from image-level representations and integrates a predefined ETF matrix as both the shared classifier and the source of class-wise query embeddings, enabling consistent feature alignment across clients (Fig. 2a). Additionally, two regularization terms (Fig. 2c, Fig. 2d) are introduced to suppress noisy negative features and enhance intra-class compactness in the latent space.
4.1 Label-Aware Disentanglement Module
In multi-label images, representing the entire image with a single pooled feature vector is suboptimal, as it compresses diverse semantics into one embedding. To address this limitation, we introduce a Label-Aware Disentanglement Module (LADM) that extracts class-specific representations from the image-level features. Our design is motivated by recent advances in multi-label learning (liu2021query2label; ridnik2023ml; xu2023label) and the conceptual similarity between multi-label classification and object detection. In object detection, multiple bounding boxes may overlap, implying that a single region can contribute to multiple object instances. Analogously, in multi-label classification, certain image regions can simultaneously provide evidence for several semantic categories.
Building upon this intuition, LADM adopts a cross-attention mechanism inspired by DETR (carion2020end). In LADM, a set of class queries attends to a grid of image tokens to extract class-specific representations. Unlike DETR, where each query corresponds to an instance and is trained with box-level supervision, LADM uses a fixed query per class and learns only from image-level labels. Each query functions as a soft region selector, aggregating spatial evidence relevant to its class while capturing contextual dependencies across regions. This design produces a set of disentangled, class-aware feature vectors that can be directly used for per-class prediction. In the FL setting, each client trains solely on its own local dataset, which may differ significantly from others due to non-IID data distributions. Such heterogeneity can cause inconsistencies in feature extraction across clients. To maintain global consistency, LADM employs a shared query matrix across all clients, ensuring that each class is represented by the same query direction. This shared design aligns feature distributions among clients and mitigates conflicts during model aggregation.
Formally, for client $k$, each sample $x_{i}^{k}\in\mathcal{D}_{k}$ ($i\in[N_{k}]$) is encoded by the backbone into a spatial feature map:
$\mathbf{F}_{i}^{k}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}},$ (3)
where $d$ denotes the channel dimension and $H\times W$ represents the spatial resolution. The feature map is then flattened into a sequence of $S=HW$ tokens:
$\mathbf{Z}_{i}^{k}=\operatorname{reshape}(\mathbf{F}_{i}^{k})^{\top}\in\mathbb{R}^{S\times d},$ (4)
and, for notational convenience, we omit the client index in subsequent equations and denote the token sequence as $\mathbf{Z}_{i}$. To preserve spatial structure, we add a fixed 2D sine–cosine positional embedding (carion2020end) to the key and value projections:
$\tilde{\mathbf{Z}}_{i}=\mathbf{Z}_{i}+\mathbf{P},\quad\mathbf{P}=\mathrm{PE}_{2\mathrm{D}}(H^{\prime},W^{\prime},d),$ (5)
where $\mathrm{PE}_{2\mathrm{D}}(\cdot)$ denotes the fixed sine–cosine positional encoding. This embedding introduces spatial awareness without adding learnable parameters, thereby ensuring a consistent inductive bias across clients.
LADM employs a 4-head cross-attention mechanism to obtain class-specific representations from the encoded image tokens. We define a shared simplex ETF matrix as
$\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C},$ (6)
where each column $\mathbf{m}_{c}$ serves both as a fixed classifier weight and as a class-specific query vector. Given the query $\mathbf{m}_{c}$ and the encoded image tokens $\tilde{\mathbf{Z}}_{i}$, LADM computes the class-specific feature using multi-head cross-attention: